OpenStudy (kittykr98):

use the function y=2x^2 -4x -6 to answer the following a) show or explain how to find the axis of symmetry b)show or explain how you would find the vertex. c)show or explain how you would find the x-intercepts. d) determine the y-intercept and explain how you would know what it is just from looking at the equation. e) graph the function

OpenStudy (kittykr98):

@jaeuni

OpenStudy (kittykr98):

y intercept is -6 i think?

OpenStudy (campbell_st):

how are you at completing the square of the function is written as \[y = 2(x^2 - 2x + ?) - 6\] complete the square inside the brackets by adding ? what do you need to add..?

OpenStudy (kittykr98):

wait a minute, you lost me?

OpenStudy (campbell_st):

do you know how to complete the square... its a method for solving quadratics..? you want a perfect square inside the brackets (x + a)^2 = x^2 + 2ax + a^2

OpenStudy (campbell_st):

or you need to get it to (x -a)^2 = x^2 - 2ax + a^2

OpenStudy (kittykr98):

this is for step A? im really sorry, its hard for me to understand things :(

OpenStudy (campbell_st):

yes it is step A... an easier option might be to use \[x = \frac{-b}{2a}\] this will give the line of symmetry and the x, value on the vertex you have b = -4 and a = 2 so substitute them to find x...

OpenStudy (kittykr98):

b = 4 a = 6 right?

OpenStudy (campbell_st):

ok... just to get the letters right... the general form is \[y = ax^2 + bx + c\] so in your equation you have a = 2 and b = -4

OpenStudy (kittykr98):

ok y=2x^2+-4x+c

OpenStudy (kittykr98):

c = 6

OpenStudy (campbell_st):

ok... can you please stop...

OpenStudy (kittykr98):

Im really confused im sorry im doing no good here

OpenStudy (campbell_st):

the equation is \[y = 2x^2 - 4x - 6\] the line of symmetry is found using \[x = \frac{-b}{2 \times a}\] you have a = 2 and b = -4 can you use the formula to calculate the value of x, which is the line of symmetry

OpenStudy (campbell_st):

so you are calculating \[x = \frac{-(-4)}{2 \times 2}\]

OpenStudy (kittykr98):

x=0 then, right?

OpenStudy (kittykr98):

-(-4) = 4 2*2 = 4 4/4 is 1

OpenStudy (campbell_st):

well you have \[x = \frac{4}{4} = 1\] so the line of symmetry is x = 1

OpenStudy (kittykr98):

i meant 1 not 0

OpenStudy (campbell_st):

ok... so part A is done Part B. The vertex is in the line of symmetry so the point is (1, ??) to find ?? just substitute x = 1 into the equation o find y

OpenStudy (campbell_st):

oops should read... the vertex is on the line of symmetry...

OpenStudy (campbell_st):

what do you get when you substitute x = 1 into the original equation..?

OpenStudy (kittykr98):

y= 2(1)^2 - 4(1) -6 y=2^2 - 4 -6 y=4 - 4 -6

OpenStudy (campbell_st):

well you substitute x = 1 so its \[y = 2\times 1^2 - 4 \times 1 - 6\] what's that value..?

OpenStudy (kittykr98):

yeah i had that but i was trying to simplify it, i guess i got ahead

OpenStudy (campbell_st):

you can type it into a calculator as its written

OpenStudy (kittykr98):

the value is -8

OpenStudy (kittykr98):

or is it -6?

OpenStudy (campbell_st):

great so that means the vertex is at (1, -8) so to find the vertex, just substitute the value for the line of symmetry into the equation and solve for y.

OpenStudy (campbell_st):

now part C let the equation = 0 to find the x intercepts \[2x^2 - 4x - 6 = 0\] divide every term by 2 \[x^2 - 2x - 3 = 0\] are you able to factor this equation..?

OpenStudy (kittykr98):

what do you mean by able to factor it?

OpenStudy (campbell_st):

well can you write the equation as the product of 2 linear factors as an example \[a^2 - 8a - 20 =(a - 10)(a + 2)\]

OpenStudy (kittykr98):

ok earlier i factored 2x+6x+4 into 2(x+2)(x+1)

OpenStudy (kittykr98):

so you mean similar to that?

OpenStudy (campbell_st):

well its a start... yest... but there are a few errors in the factoring... 2 x -1 = -2 you need -3 adding them 2 + -1 = 1 you want -2... so you need a little rethinking

OpenStudy (campbell_st):

of you factor out the 2 you get \[2(x^2 - 2x - 3) = 0\]

OpenStudy (kittykr98):

ok so im trying to understand

OpenStudy (campbell_st):

so find the factors of -3 that add to -2 the larger factor is negative and the smaller factor is positive

OpenStudy (kittykr98):

ok hold on

OpenStudy (kittykr98):

when multiplied it needs to equal -3 but when added it needs to equal -2, right?

OpenStudy (campbell_st):

thats correct

OpenStudy (kittykr98):

-3 and positive 1

OpenStudy (campbell_st):

great so now you can write the equation as \[2(x -3)(x+1) = 0\] now if either of the linear factors x -3 or x + 1 equal zero, the whole thing will equal zero. so you need to find the x values that make x - 3 = 0 and x + 1 = 0 this will be where the parabola cuts the x-axis

OpenStudy (kittykr98):

3-3=0 and -1+1=0

OpenStudy (campbell_st):

great... so the x-intercepts are at (-1, 0) and (3, 0) those points come from the solutions you just game me. Now to Part D this is probably the easiest... substitute x = 0 into the original equation to find y.... this will be the y-intercept.

OpenStudy (campbell_st):

any thoughts on the y-intercept..?

OpenStudy (kittykr98):

I know that it is -6 but do not know how to explain why

OpenStudy (kittykr98):

Y=2(0)^2 - 4(0) -6 Y= 4 - 4 - 6 Y= 6 So the y intercept is 6, right?

OpenStudy (campbell_st):

ok... well replace x with 0 and calculate \[y = 2 \times 0^2 - 4 \times 0 - 6\] so almost done... now Part E plot the points on the number plane Vertex (1, -8) x- intercepts (-1, 0) and (3, 0) y- intercept (0, -6) draw a line through all 4 points and you get a parabola. you can use this site to check https://www.desmos.com/calculator hope it all helped.

OpenStudy (kittykr98):

Sorry I had to switch to the mobile app because my laptop died and it's all crazy. You've been a great help! And now finally I can go to bed! Thank you so much!! Can I tag you in a question tomorrow?

OpenStudy (campbell_st):